# Stretching Beyond the Obvious: A Gradient-Free Framework to Unveil the Hidden Landscape of Visual Invariance

### v1

@@ 5 Discussion
Finally, the model-agnostic and gradientless nature of SnS positions it as a powerful new tool for neuroscience, particularly for systems where high-fidelity “digital twins” may be hard to develop. Our demonstration of SnS’s efficacy with simulated sparse neural recordings directly addresses a critical experimental constraint, paving the way for in vivo applications. This could lead to the discovery of new, hierarchical invariances in biological visual systems, extending our understanding of visual object representations in both primates [20, 30] and other species [42, 43, 44, 28, 45, 46, 47].
Limitations and broader impacts. While SnS has proven effective in uncovering novel, meaningful invariances, its current evolutionary, model-agnostic implementation presents opportunities for refinement and broader application. One avenue is to develop hybrid SnS variants that incorporate gradient-based optimization where network differentiability is available. Such an approach could potentially accelerate convergence or fine-tune the search, leveraging system-specific knowledge while retaining model-agnostic capabilities for biological or other black-box targets. The hierarchical invariance characterization we introduced could be significantly expanded. Future work should apply SnS across a more diverse array of subject networks used for generation and “observer” networks for evaluation, extending beyond CNNs to architectures with different inductive biases, such as Vision Transformers. This would help determine the generality of the observed hierarchical patterns and how architectural choices influence learned invariances. Furthermore, our finding that invariant images of standard ResNet50 are better understood by larger, high-performing standard observer networks warrants deeper investigation. Unraveling the mechanisms behind this phenomenon could shed light on the complex interplay between model scale, feature learning, and generalization. Finally, while SnS’ robustness to significant representational subsampling provides a strong foundation for its neuroscientific applicability, the crucial next step is direct validation in biological systems. Future research should focus on adapting SnS for in vivo experiments, addressing challenges such as neural signal variability, and optimizing stimulus presentation within experimental constraints to truly bridge the gap between computational characterization and biological understanding of invariance.
@@ A.1 Pseudocode for the SnS algorithm
Here we present the pseudo-code (Algorithm 1) for the SnS algorithm. We used library implementations for the CMA-ES gradient-free optimizer [48] and the Pareto front computation [49]. Importantly, we introduce the following should_stop criteria for our experiments.
@@
$\phi$triangleright Target network
$\psi$triangleright Image generator
$T\geq 0$triangleright Maximum number of iterations
@@
$\bm{x}_{\mathrm{ref}}\in\mathbb{R}^{C\times H\times W}$triangleright Common choice is $\bm{x}_{\mathrm{ref}}=\bm{x}_{\star}$
@@
     $\bm{x}_{t}=\psi\left(\bm{\xi}_{t}\right)$triangleright Compute new candidate stimuli
     $\mathbf{a}^{\kappa,\ell}=\Gamma\left(\bm{x}_{t},\phi^{\kappa,\ell}\right)$triangleright Measure activations in target layers $\kappa,\ell$
@@
     $\bm{\xi}_{t+1}\leftarrow\mathtt{evolve}\left(\mathcal{P},\bm{\xi}_{t}\right)$triangleright Evolution strategy CMA-ES optimization
@@ A.2 Experimental Hyperparameters
The generative model $\psi$ is instantiated as a pretrained deep neural network variant, specifically the fc7 configuration from [32].
@@
A.3 Optmization Convergence
@@
Figure S1: SnS images fall into functionally relevant activation regimes. Distributions of readout target unit activations (normalized by peak natural stimulus response) for SnS-generated images: invariant (stretched at pixel, mid-, high-levels) and adversarial (pixel-space). Each distribution pulls together all optimizations performed in that specific condition (n = 77 target neurons x 10 random seeds = 770 experiments per condition). Comparison with natural image activations (same/different category) confirms that SnS invariant images elicit strong target responses, while adversarial examples achieve significant suppression.
@@
Although some optimizations terminated at maximum iterations (adversarial: 97.92% (robust), 93.12% (standard); invariant: 63.38%, 35.71%, 21.30% (robust: low_level, mid_level, high_level); 86.75%, 32.60%, 5.84% (standard: low_level, mid_level, high_level)) , final populations consistently reached functionally relevant activation regimes. Most invariant images achieved activations within the top 0.1% of natural image responses, being comparable or more extreme than the readout activation of images of the same category encoded by the readout target unit (Fig. S1). Adversarial examples drove activations to the lowest $1^{st}$ percentile of the natural images, thus significantly suppressing target unit activation w.r.t. average activation with a natural image.
A.4 Computational Resources
@@
The SnS source code is openly available in the GitHub repository at https://github.com/zoccolan-lab/SnS. Moreover the experimental data is accessible via [50].
Appendix B Additional details regarding human subjects experiment
@@
The experiment posed minimal risks to participants. It is unlikely but conceivable that the brief presentation of images and masks could be hazardous to a subset of individuals with photosensitive epilepsy: out of an abundance of caution, a warning was shown in the description of the task on the Prolific website to be viewed by participants before joining the study. The title and description of the task on the Prolific website are reproduced as follows:
@@
Please see Fig. S2 for the additional in-task instructions provided to participants. The task was implemented using the JsPsych library [51] and the JsPsychPsychophysics plugin [52].

### v2

@@ 5 Discussion
Finally, the model-agnostic and gradientless nature of SnS positions it as a powerful new tool for neuroscience, particularly for systems where high-fidelity “digital twins” may be hard to develop. Our demonstration of SnS’s efficacy with simulated sparse neural recordings directly addresses a critical experimental constraint, paving the way for in vivo applications. This could lead to the discovery of new, hierarchical invariances in biological visual systems, extending our understanding of visual object representations in both primates (DiCarlo et al., 2012; Ponce et al., 2019) and other species (Muratore et al., 2025; 2022b; Tafazoli et al., 2017; Tong et al., 2023; Soto and Wasserman, 2016; Wood and Wood, 2015; Schluessel et al., 2012).
Limitations and broader impacts. While SnS has proven effective in uncovering novel, meaningful invariances, it presents opportunities for refinement and broader application. One avenue is to use SnS for further, more extensive geometric characterizations of the invariance manifolds, expanding the work presented in Supplementary Material, Section H. Another interesting direction is to use alternative generative models, such as large-scale GANs and diffusion models, to search the image space for invariant images. However, these models impose very strong priors on the structure of the synthesized images, which risks evolving exceedingly complex photorealistic details that are unlikely to be encoded by units in low- and mid-level processing stages (Wang and Ponce, 2024). Yet another possibility is to apply SnS to characterize invariance in other modalities, such as audition, using a generator capable of producing audio waveforms or spectrograms from a low-dimensional latent code - e.g., SpecGAN, WaveGAN (Donahue et al., 2018). Returning to the visual domain, our work has demonstrated SnS’ generative capabilities using eight different neural models (Fig. 4 and S9), belonging to three different architectural families, with parameter counts spanning roughly an order of magnitude (from 11.7M for ResNet-18 to 86.6M for ViT). This provides a strong test of our framework’s generalizability, although future work could test SnS with even larger models (e.g., DINO (Caron et al., 2021), ViT22-B (Dehghani et al., 2023)). Finally, by pinpointing where the perceptual misalignement between robustified networks and humans is strongest (i.e., in higher-order invariances), SnS could guide the development of models that achieve more human-aligned invariance. For instance, one could design a visual diet where “good” (i.e., human-like) and “bad” (i.e., nonhuman-like) invariant images are explicitly used in the training of the network.
In terms of application to neuroscience studies, while SnS’ robustness to significant representational subsampling provides a strong foundation for its neuroscientific applicability, the next step is direct validation in biological systems. Future research should focus on adapting SnS for in vivo experiments, addressing issues such as neural signal variability, the limited duration of neuronal recordings, and the need to optimize stimulus presentation in closed loop experiments, similarly to what was already achieved with XDREAM (Ponce et al., 2019). One challenge will be to leverage the wealth of neuronal data afforded by high-density silicon probes (Siegle et al., 2021), that can simultaneously record from hundreds to thousands of units. One strategy would be to apply SnS in parallel to multiple recorded neurons with non-overlapping receptive fields (RFs), thus independently evolving the distinct portions of the image processed by each unit. Conversely, for neurons with overlapping RFs, one could first screen them using natural images (Ponce et al., 2019) to search for subpopulations of units with similar selectivity. SnS could then be applied to the subpopulation to find the collective invariance field of the neuronal ensemble.
6 Reproducibility statement
To enhance the reproducibility of our research, we provide the following resources. The SnS algorithm is described in detail within the main paper’s Methods Section 3.1, with its corresponding pseudocode available in the Supplementary Material (see Algorithm 1). Furthermore, the complete source code used to conduct the experiments is available at https://github.com/zoccolan-lab/SnS, while experimental data were open sourced at Tausani et al. (2025). For clarity on our experimental setup, a general description of the experimental parameters can be found in the methods section. For a more granular understanding, further information on methodological details, hyperparameter settings and the computational resources used is reported in Supplementary Material, Section A. The Supplementary Material also includes a comprehensive section on the statistical analysis employed in our experiments (see Section B.1).
7 Ethics statement
The human behavioral study was conducted under a pre-existing, approved Institutional Review Board (IRB) protocol. The task, which involved participants viewing images on a computer screen and making responses by clicking buttons, presented no more than minimal risk. Further details, including a complete reproduction of test instructions and information on participant compensation, are provided in Supplementary Material Section B.
@@ A.1 Pseudocode for the SnS algorithm
Here we present the pseudo-code (Algorithm 1) for the SnS algorithm. We used library implementations for the CMA-ES gradient-free optimizer (Hansen et al., 2019) and the Pareto front computation (Fortin et al., 2012). Importantly, we introduce the following should_stop criteria for our experiments.
@@
In terms of computational complexity, both generator-based image synthesis from latent codes and the extraction of activations from the target network require only a single feedforward pass per iteration, whose computational cost, while depending on network size, is negligible in practice for all the architectures examined in our study, which span roughly an order of magnitude of parameter counts (from 11.7M for ResNet-18 to 86.6M for ViT). Consequently, the dominant computational bottleneck is the CMA-ES optimizer, whose running time scales quadratically with the dimensionality $d$ of the search space. In practice, however, this does not pose a limitation: CMA-ES has been shown to be a highly effective method for gradient-free optimization of neural activity in vivo (Wang and Ponce, 2022a), and it remains computationally efficient under our hardware configuration as well (see Section A.6).
@@
$\phi$ $\triangleright$ Target network
$\psi$ $\triangleright$ Image generator
$T\geq 0$ $\triangleright$ Maximum number of iterations
@@
$\bm{x}_{\mathrm{ref}}\in\mathbb{R}^{C\times H\times W}$ $\triangleright$ Common choice is $\bm{x}_{\mathrm{ref}}=\bm{x}_{\star}$
@@
  $\bm{x}_{t}=\psi\left(\bm{\xi}_{t}\right)$ $\triangleright$ Compute new candidate stimuli
  $\mathbf{a}^{\kappa,\ell}=\Gamma\left(\bm{x}_{t},\phi^{\kappa,\ell}\right)$ $\triangleright$ Measure activations in target layers $\kappa,\ell$
@@
  $\bm{\xi}_{t+1}\leftarrow\mathtt{evolve}\left(\mathcal{P},\bm{\xi}_{t}\right)$ $\triangleright$ Evolution strategy CMA-ES optimization
@@ A.2 Experimental Hyperparameters
The generative model $\psi$ is instantiated as a pretrained deep neural network variant, specifically the fc7 configuration from Dosovitskiy and Brox (2016).
@@
Figure S1: Examples of solutions in the final Pareto Front. Visualization of the final Pareto fronts for $n=4$ example units in Invariance experiments (top) and $n=4$ example units in Adversarial Attack experiments (bottom). In all reported experiments, targets were located in the readout layer, while pixel space was used as lower representation space. The ordinate axis represents the activation reduction of a readout unit of a robust or standard ResNet50 with respect to the activation produced by a reference image (MEI), when SnS is applied to synthesize either adversarial or invariant images, while the abscissa represents the $L_{2}$ pixel distance between these synthesized images and the reference. In the Invariance Experiments, SnS maintains a substantial fraction of the reference image activation (regardless of whether the optimization fully converges) while allowing the synthesized images to diverge markedly from the reference in pixel space. Conversely, in the Adversarial Attack setting, SnS reliably produces images that almost completely suppress the response of the target unit, yet remaining comparatively close to the MEI in the pixel space representation. Note the different scale range of the abscissa and ordinate axes between the top and bottom plots.
A.3 Pareto front computation
Multi-objective optimization often involves conflicting goals (e.g., stretching vs. squeezing, see Fig. 1). One way of dealing with such problems is by structuring the fitness function as a linear combination of the different optimization objectives (weighted-sum methods). However, these methods often require fine parameter tuning to prevent one objective from dominating. Unlike weighted-sum methods, Pareto fronts (PFs) handle trade-offs without parameter tuning. At each optimization iteration, candidate solutions (n=50) are ranked iteratively into PFs in the two-objective space. Each PF contains mutually non-dominant solutions (i.e. one is not better than another in both objectives). To get a full ordering, solutions are sorted within each front:
Invariance Experiments: In this scenario, solutions on the same Pareto front were prioritized based on their proximity to the target unit’s reference activation level ($\min_{x}\parallel\Gamma(\bm{x},\phi^{\ell})-\bm{x}^{\ell}_{\mathrm{ref}}\parallel_{2}$). This exploitative strategy was motivated by the empirical difficulty of converging towards the target activation level while at the same time maximizing input distance.
Adversarial Attacks: Solutions on the same Pareto front were selected for the next generation with uniform probability (random ordering). This promotes exploration along the front, preventing premature convergence to a single type of adversarial strategy.
CMA-ES then updates its distribution using the best $30\%$ of this ranking. At the end of the optimization, all solutions are used to compute a global PF, which is thus fully data-driven (see Fig. S1 for several examples). In our study, we analyzed only the last solution from this final PF for each experiment.
A.4 Separability of the invariant images in the pixel representation
To assess whether the representation spaces where the stretching was applied yielded sets of invariant images that were distinguishable at the pixel level, we performed an image classification analysis. We generated invariant images for $n=77$ readout layer units (i.e., 77 ImageNet classes) by stretching the representations in three different processing stages: (i) low_level (pixel space), (ii) mid_level, and (iii) high_level. For each unit and each representation space condition, we performed $10$ independent optimization runs with different random seeds, yielding a total dataset of $77\times 3\times 10=2310$ images. Principal Component Analysis (PCA) was applied to the raw pixel representations of all images. The first $k$ principal components were then used as input features to train a Support Vector Classifier with a radial basis function kernel to discriminate the class to which the invariant images belonged (i.e., low_level, mid_level, or high_level). 5-fold cross-validation was used.
A.5 Optimization Convergence
A.5.1 SnS images fall into functionally relevant activation regimes
@@
Figure S2: SnS images fall into functionally relevant activation regimes. (a) Distributions of readout target unit activations (normalized by peak natural stimulus response) for SnS-generated images: invariant (stretched at pixel, mid-, high-levels) and adversarial (pixel-space). Each distribution pulls together all optimizations performed in that specific condition (n = 77 target neurons x 10 random seeds = 770 experiments per condition). Comparison with natural image activations (same/different category) confirms that SnS invariant images elicit strong target responses, while adversarial examples achieve significant suppression. (b) The average convergence curves ($\pm$ SD) for the experiments illustrated in Fig. 2 (invariance and adversarial attacks, n = 770 runs per condition). Both optimization objectives (i.e., activation of the target units and Euclidean distance from the MEI in pixel space) are shown as a function of iteration number. Both objectives converged to stable asymptotic values consistent with the intended optimization target (i.e., the search for invariant or adversarial images). In particular, the Euclidean distance reached a plateau, indicating that the images no longer continued to evolve significantly following an initial period of convergence.
Table S1: Mean activations of target units to SnS-synthesized images from optimizations that failed to converge. Reported activations are normalized by the response to the maximally activating natural image for the corresponding unit. N denotes the number of runs in which the early-stopping criterion was not reached for each experimental condition, out of a total of 770 optimization runs per condition. Successful maintenance of activation levels in Invariance runs would be indicated by positive activation values in the Mean column near to or exceeding 1, while successful suppression of activity in Adversarial runs would result in low or negative values.
Experiment Layer Model Type Mean SEM N
Invariance low_level Standard 0.7682 0.0103 668
Invariance mid_level Standard 0.9581 0.0110 251
Invariance high_level Standard 0.9623 0.0233 45
Adversarial low_level Standard -0.1920 0.0027 718
Invariance low_level Robust $L_{2}$ 0.8467 0.0056 489
Invariance mid_level Robust $L_{2}$ 0.9104 0.0059 276
Invariance high_level Robust $L_{2}$ 0.9271 0.0075 166
Adversarial low_level Robust $L_{2}$ -0.1547 0.0036 755
@@
Although some optimizations terminated at maximum iterations (adversarial: 97.92% (robust), 93.12% (standard); invariant: 63.38%, 35.71%, 21.30% (robust: low_level, mid_level, high_level); 86.75%, 32.60%, 5.84% (standard: low_level, mid_level, high_level)), final populations consistently reached functionally relevant activation regimes. Most invariant images achieved activations within the top 0.1% of natural image responses, being comparable or more extreme than the readout activation of images of the same category encoded by the readout target unit (Fig. S2a). Adversarial examples drove activations to the lowest $1^{st}$ percentile of the natural images, thus significantly suppressing target unit activation with respect to average activation with a natural image. As shown in Table S1, this was true also when considering only those units for which the early-stopping criterion was not reached, i.e., the optimizations that terminated at the maximum number of iterations (500). Moreover, as illustrated by the convergence curves in Fig. S2b, the average trajectories of both optimization objectives (activation of the target readout units and Euclidean distance from the MEI in pixelspace) rapidly approached stable asymptotic values consistent with the respective optimization goals (i.e., invariant and adversarial). In particular, the Euclidean distance plateaued, indicating that the optimized images no longer underwent substantial changes after a certain number of iterations.
A.5.2 Invariant solutions are maximally distinct from their references and heterogeneous across multiple runs
Figure S3: Representation of the invariant images across the feedforward hierarchy. (a) Average distance between the invariant images generated by SnS and their reference MEIs across the different stages of a robust ResNet50. Euclidean distances were normalized by the mean, within-category distance of Imagenet images. Different lines indicate the different stretching layers used in the optimization (also reported in the $x$ axis). The average distance between randomly selected natural images from Imagenet (solid gray line) and the average distance between multiple MEIs generated via XDREAM are reported for comparison. (b) Same as (a) for standardly trained network.(c) Normalized average distance between the invariant images generated by SnS across the different stages of a robust ResNet50. Different lines indicate the different stretching layers used in the optimization. The average distance between randomly selected natural images from Imagenet (solid gray line) and the average distance between multiple MEIs generated via XDREAM are reported for comparison. (d) Same as (c) for standardly trained network.
To quantify how much the invariant images generated by SnS (when targeting a given processing stage) diverged from the reference MEI ($\bm{x}^{\star}$) across the depth of ResNet50, we computed the Euclidean distance between their activation vectors at each layer of the network (Representational Distance Analysis in short). Again, 10 different invariant images were produced for $n=77$ distinct readout units by stretching representations at layers low_level, mid_level, and high_level. We calculated the mean ($\pm$ SEM) Euclidean distance between each of these $77\times 10$ invariant images and the corresponding reference images at every convolutional, pooling and linear layer of the network, separately for each stretching stage. Analogously, to quantify heterogeneity between different invariant solutions, for each aforementioned stage we measured the average pairwise distance between invariant images, then averaged across each target neuron ($n=77$). These distances were compared against three control metrics:
In our analysis (see Fig. S3) we normalized all Euclidean distances by the mean, within-category distance of Imagenet images. Fig. S3a-b show that invariant images that were generated to be maximally different from the MEI in a specific layer had the largest normalized distance in that layer, as compared to invariant images obtained by applying the stretching in other layers. Moreover, this distance consistently surpassed the one among natural images (both within the same category and randomly sampled) and multiple MEIs of the same unit, demonstrating SnS’s efficacy in exploring larger spans of a unit’s invariance manifold, as compared to searching multiple times for different MEIs. Moreover, Fig. S3c-d show that distances between invariant images were significantly above 0 for all stretching layers (pixel space, mid, and late layers) in both standard and robust ResNet50, indicating a substantial diversity between invariant images. Heterogeneity is lowest in the pixel space (standard: 0.69, robust: 1.03) and higher in the mid (standard: 1.24, robust: 1.66) and late layers (standard: 0.85, robust: 1.27). Interestingly, invariances in the mid-layers of both networks - and those in the late layer of the robust network - are more heterogeneous than both random ImageNet pairs and multiple MEIs for the same unit. Instead, the invariant images in the late layer of the standard network exceed MEI heterogeneity, but not that of random ImageNet pairs. This heterogeneity is supported by SnS random initialization, which promotes broad exploration of the invariance landscape across multiple runs for the same unit and reduces the likelihood of converging to the same local minimum.
A.6 Computational Resources
@@
On this setup, a full optimization of a single SnS run (500 iterations) lasted approximately 120 seconds. The complete source code used to conduct the experiments is available at https://github.com/zoccolan-lab/SnS. Moreover the experimental data is accessible via Tausani et al. (2025).
Appendix B Procedure to probe the interpretability of the invariant images by humans and observer networks
For our interpretability experiments, we chose $12$ ImageNet categories (goldfish, chickadee, frilled lizard, Dungeness crab, Staffordshire bull terrier, cicada, candle, grand piano, minibus, reflex camera, soccer ball, cup). For each category, we selected 10 reference images $\bm{x}_{\mathrm{ref}}$ and, for each of them, we generated one invariant image by stretching its representation in low_level, mid_level, and high_level in both standard and robust ResNet50. This yielded $6$ kinds of invariant images per reference image, where invariance was always defined relative to the readout unit corresponding to the reference image’s category. For each category, we also included in the classification pool the 10 reference images and 20 MEIs (generated with XDREAM) of the readout unit corresponding to the chosen category (10 for the standard and 10 for the robust network). This yielded a total of 12 categories $\times$ 10 randomly sampled images $\times$ 9 stimulus types (i.e., 1 reference image, 2 MEIs and 6 invariant images). Observer networks had to classify all $1080$ of these images; humans had to classify a subset of $540$ images (i.e., $5$ rather than $10$ randomly sampled images per category). Classification accuracy (12-AFC percent correct) was calculated for each of the 9 stimulus types. To assess the generalizability of our findings, we repeated the same experiment with invariant images generated by two additional architectures (ResNet18 and VGG16_bn, in both standard and robust versions; see details above). The pool of observer networks was identical to the previous experiment. As for ResNet50, 3 levels of stretching were applied: Pixel space, mid level (VGG16_bn: 2nd convolutional layer in the 3rd block; ResNet-18: 4th convolutional stage in layer 2) and high level (VGG16_bn: 3rd convolutional layer in the 5th block; ResNet-18: 4th convolutional stage in layer 4).
The psychophysics experiment administered to human participants was conducted as follows. After clicking on a fixation cross at the start of each trial, participants were presented with an image for 50 ms at $\sim$10 of visual angle. After a 20ms blank screen, a backward mask image (random RGB pixel noise) was then presented for 200ms. Images were presented in random order. Participants then clicked one of 12 category buttons that subsequently appeared in a circular arrangement, with a 10-second timeout (Fig. 4a). To avoid directional bias, the sequence of category buttons was randomly rotated for each trial.
@@ Appendix B Procedure to probe the interpretability of the invariant images by humans and observer networks
The experiment posed minimal risks to participants. It is unlikely but conceivable that the brief presentation of images and masks could be hazardous to a subset of individuals with photosensitive epilepsy: out of an abundance of caution, a warning was shown in the description of the task on the Prolific website to be viewed by participants before joining the study. To ensure consistent stimulus presentation, participants were required to complete the task on a desktop or laptop computer; access via tablets and smartphones was programmatically blocked. The title and description of the task on the Prolific website are reproduced as follows:
@@
Table S2: Statistical results from omnibus test (type II Wald chi-square) and post-hoc pairwise comparisons to test differences among stimulus types in the human image classification experiments. All $p$ values are adjusted using the Benjamini-Hochberg correction, with significant values in bold.
Human subjects
Comparison z-value p-value
Omnibus test (effect of stimulus type) - <0.0001
Average Robust Stretch vs Average Standard Stretch 36.63 <0.0001
Robust layer4_conv7 vs Robust layer3_conv1 -4.94 <0.0001
Standard layer4_conv7 vs Standard layer3_conv1 -8.41 <0.0001
Robust layer4_conv7 vs Robust pixel space -9.52 <0.0001
Standard layer4_conv7 vs Standard pixel space 7.49 <0.0001
Robust layer3_conv1 vs Robust pixel space -4.66 <0.0001
Standard layer3_conv1 vs Standard pixel space -0.99 0.32
Robust MEI vs Standard MEI 32.96 <0.0001
Please see Fig. S4 for the additional in-task instructions provided to participants. The task was implemented using the JsPsych library (De Leeuw, 2015) and the JsPsychPsychophysics plugin (Kuroki, 2021).